OpenStudy (anonymous):

what is 1/5+3/10? it is not necessary to simplify answer. please help i dont get it and explain

OpenStudy (anonymous):

5/10 OR 1/2

OpenStudy (anonymous):

\[\frac{1}{5}=\frac{2}{10}\] and so \[\frac{1}{5}+\frac{3}{10}=\frac{2}{10}+\frac{3}{10}=\frac{5}{10}\]